Faz, Inc., manufactures and sells two products: Product X0 and Product W7. Data concerning the expected production of each product and the expected total direct labor-hours (DLHs) required to produce that output appear below:

Expected Production Direct Labor-Hours Per Unit Total Direct Labor-Hours
Product X0 1,500 4 6,000
Product W7 420 2 840
Total direct labor-hours 6,840

The direct labor rate is $21.60 per DLH. The direct materials cost per unit is $146.50 for Product X0 and $128 for Product W7.

The company is considering adopting an activity-based costing system with the following activity cost pools, activity measures, and expected activity:

Estimated Expected Activity
Activity Cost Pools Activity Measures Overhead Cost Product X0 Product W7 Total
Labor-related DLHs $ 274,078 6,000 840 6,840
Production orders orders 19,148 520 720 1,240
Order size MHs 238,134 4,020 4,120 8,140
$ 531,360

The overhead applied to each unit of Product X0 under activity-based costing is closest to: (Round your intermediate calculations to 2 decimal places.)

rev: 03_25_2018_QC_CS-119201

Multiple Choice

  • $310.74 per unit

  • $249.04 per unit

  • $78.39 per unit

  • $244.02 per unit
Answer #1

calculate overhead to each unit

Product X0
Labor related 274078/6840*6000 = 240419
Production orders 19148/1240*520 = 8030
Order size 238134/8140*4020 = 117604
Total 366053
Unit 1500
Overhead cost per unit 244.02

So answer is d) $244.02 per unit